# Towards Multi-Modal Animal Pose Estimation: An In-Depth Analysis

### v1

@@ 4.1.1. Single Animal
2D single-frame single-animal pose estimation aims to accurately estimate the 2D pose of an animal from a single image. These methodologies share a conceptual foundation with HPE (Zheng et al., 2020) and can be broadly classified into two categories: heatmap-based thods. The former predicts the locations of an animal’s body joints using heatmaps for supervision, necessitating a subsequent postprocessing step to identify the final keypoints based on the predicted heatmaps. In contrast, the latter approach directly regresses the keypoint coordinates of the animal’s body joints, employing an end-to-end deep neural network to establish a mapping from the input image to the body keypoints. LABEL:fig:single_animal illustrates the general architecture of 2D single-frame single-animal pose estimation methods.
@@
Heatmap-based frameworks such as (Pereira et al., 2019; Kim et al., 2022; Li and Lee, 2023; Li et al., 2020a; Li and Lee, 2021b; Graving et al., 2019; Li et al., 2023; Mathis et al., 2018) have become popular in 2D APE. These approaches are structured around creating a 2D heatmap, interpreted as a probabilistic representation of body part locations within the image. The architecture of a typical heatmap-based scheme includes i) a CNN backbone to generate heatmaps from input images, followed by ii) a post-processing step to finalize the pose estimations based on these predictions (LABEL:heatheat). There are some common architectures to input an image and output the corresponding heatmaps in step i); for instance, (Kim et al., 2022; Li and Lee, 2021b; Li et al., 2023; Mathis et al., 2018) uses the ResNet (He et al., 2016) and its variants to produce the heatmap; (Li and Lee, 2023) adapts a variant of the HRNet (Sun et al., 2019) as the backbone; (Li et al., 2020a) applies the stacked hourglass (Newell et al., 2016) architectures. Subsequently, in step ii), the predicted heatmaps are then processed to produce precise 2D spatial coordinates for each body joint using techniques such as peak detection (Graving et al., 2019; Mathis et al., 2018; Pereira et al., 2019). Although heatmap-based approaches are well known for their robustness against occlusions and variations in appearance, their computational complexity has been critiqued for impeding end-to-end training. Nonetheless, DeepPoseKit (Graving et al., 2019) proposes an efficient end-to-end heatmap-based approach.
@@
Regression-based approaches directly regress the coordinates of joint locations as continuous values, employing CNNs to learn a mapping from the input image to the joint coordinates. The output layer of the network consists of multiple units, each representing the $x$ and $y$ coordinates of a joint, and is trained to minimize the distance between the predicted and ground truth coordinates. Although regression-based approaches are less common than heatmap-based methods, they have shown great promise (Lu and Koniusz, 2022). One notable work (Lu and Koniusz, 2022) adopts a regression-based approach to estimate keypoint positions directly, bypassing the intermediate heatmap representation (LABEL:RegressionRegression). Their Few-shot Keypoint Detection (FSKD) model is designed to detect keypoints on unseen species with minimal annotated data, which is particularly useful in challenging scenarios where the species have not been previously encountered or the available annotated data is limited. Regression-based approaches offer the advantage of direct joint location estimation without additional peak detection steps, which can benefit real-time inference. However, they may face long-range 2D offset regression challenges, which may lead to a notable performance gap compared to heatmap-based methods.
@@
Table 5. RGB-based 2D APE Methods. In the column ”I,V”: ”I,V” is the methods that can perform pose estimation on both an image and videos, whereas, ”I” is the APE methods that do only on an image. Column ”L” indicates the learning types as follows: ”S” is supervised methods, ”U” is unsupervised methods, ”SF” is self-supervised methods, ”SM” is semi-supervised methods. [Uncaptioned image] is CNN architecture, [Uncaptioned image] is the use of the attention mechanism, [Uncaptioned image] is GAN, [Uncaptioned image] is underlying architecture as vision transformer, [Uncaptioned image] is Multi-Scale Domain Adaptation Module (MDAM). [Uncaptioned image] is the pose estimation of the candidate algorithm given in the keypoint pose format. [Uncaptioned image] is single-camera-view, [Uncaptioned image] is multi-camera-views. [Uncaptioned image] is ”indoor” experimental set-up and [Uncaptioned image] is ”in-the-wild” experimental set-up. Symbols inside the {} indicate that both categories are used in that method.
@@
1 2018 DeepLabCut (Mathis et al., 2018) I,V S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Mammals RMSE (pixels)
2 2019 LEAP (Pereira et al., 2019) I,V S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Fly, Mouse RMSE (pixels)
3 2019 DeepPoseKit (Graving et al., 2019) I,V S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], {[Uncaptioned image], [Uncaptioned image]} Zebra (Zuffi et al., 2019b) ([Uncaptioned image]), Euclidean Error,
Desert locust, Fly ([Uncaptioned image]) Bayesian Inference (Graving et al., 2019)
4 2020 Li et al. (Li et al., 2020a) I S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Fly, Worm, RMSE, PCK, AUC (pixels),
Zebrafish PI-AUC (4-45 pixels),
PI-PCK (5&15 pixels),
PI-RMSE (pixels)
5 2021 Synthetic-to-Real (Li and Lee, 2021b) I U [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Horse, TigDog(Del Pero et al., 2017) PCK@0.05 (%)
6 2022 Kim et al. (Kim et al., 2022) I SM,U [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] {[Uncaptioned image], [Uncaptioned image]} TigDog(Del Pero et al., 2017), PCK@0.05 (%)
Synthetic Animal (Mu et al., 2020),
Animal Pose (Cao et al., 2019)
7 2023 ScarceNet (Li and Lee, 2023) I S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] AP-10K (Yu et al., 2021b), mAP (%), PCK@0.05 (%)
TigDog(Del Pero et al., 2017)
8 2023 D-Gen (Li et al., 2023) I S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] AP-10K (Yu et al., 2021b), OKS, PCK@0.05 (%)
Animal Pose (Cao et al., 2019),
Animal Kingdom (Ng et al., 2022a)
@@
9 2022 FSKD (Lu and Koniusz, 2022) I S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], {[Uncaptioned image], [Uncaptioned image]} Animal Pose (Cao et al., 2019) PCK
@@
10 2022 SIPEC (Marks et al., 2022) I,V S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Primates, Mice Pearson Correlation,
RMSE (pixels), mAP (%),
IoU, Dice Coefficient
11 2022 DepthFormer (Liu et al., 2022b) I S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] AP-10K (Yu et al., 2021b) OKS
12 2024 Straka et al. (Straka et al., 2024) I S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Lynx PCK, OKS, AP
13 2024 SuperAnimal-Quadruped I,V S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] AP-10K (Yu et al., 2021b), RMSE (pixels), mAP,
(Shaokai Ye and Mathis, 2024) Animal Pose (Cao et al., 2019) Normalized Error
@@
14 2017 OpenPose (Cao et al., 2017) I,V S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], {[Uncaptioned image], [Uncaptioned image]} COCO (Lin et al., 2014b), MPII (Andriluka et al., 2014) mAP (%)
15 2021 Negrete et al. (Blanco Negrete et al., 2021) I S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] MacaquePose (Labuguen et al., 2021), AP
MSCOCO (Lin et al., 2014a)
16 2021 Farahnakian et al. (Farahnakian et al., 2021) I,V S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Pig(Nath et al., 2019) Error (pixels), RMSE (pixels)
@@
17 2022 DLCRNet, maDLC (Lauer et al., 2022) I,V S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], {[Uncaptioned image], [Uncaptioned image]} 3 Mice ([Uncaptioned image]), 1 Mice ([Uncaptioned image]), RMSE (pixels),
2 marmosets ([Uncaptioned image]) mAP, PCK
Bottom-up & Top-Down:
@@
18 2022 SLEAP (Pereira et al., 2022) I,V S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Flies, Bees, mAP (%), PCK,
Mice, Gerbils OKS, Localization
@@
19 2024 SuperAnimal-TopViewMouse I,V S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] AP-10K (Yu et al., 2021b), RMSE (pixels), mAP,
(Shaokai Ye and Mathis, 2024) Animal Pose (Cao et al., 2019) Normalized Error
4.1.2. Multiple Animals
@@
We have seen significant developments in markless pose estimation in individual animals. However, extending these methodologies to multiple animals poses unique challenges, particularly in studying social behaviours or animals in their natural environments (Yang et al., 2022b; Gong et al., 2022). For example, it is common for the body parts of one animal to be obscured by another in a multi-animal pose estimation. Another challenge arises from scale variation, where animals at different distances from the camera show differences in apparent sizes. Inter-species variation also introduces additional complexity due to differences in body structures and poses. However, given these challenges, 2D multiple-animal pose estimation research has still garnered significant attention due to its potential applications in various domains, such as animal behaviour analysis, wildlife monitoring, and ecological studies. Methodologies developed in this field are categorized into two main categories: top-down methods (Marks et al., 2022; Liu et al., 2022b; Straka et al., 2024; Shaokai Ye and Mathis, 2024) and bottom-up methods (Cao et al., 2017; Blanco Negrete et al., 2021; Farahnakian et al., 2021; Lauer et al., 2022) (LABEL:fig:multi_animal).
@@
Top-down approaches for multiple-animal pose estimation (LABEL:aaa) generally involve a two-step process: first i) detecting each animal in the input image using pre-trained object detectors and then ii) applying 2D single-animal pose estimation techniques to estimate the pose for each detected instance. Common object detectors in step i) include Faster R-CNN (Ren et al., 2015), the Single Shot MultiBox Detector (SSD) (Liu et al., 2016), You Only Look Once (YOLO) v4 (Gong et al., 2022), and RetinaNet (Lin et al., 2017); common 2D APE methods in step ii) includes variants of ResNet, HRNets and transformer. In the early work, SIPEC (Marks et al., 2022) uses SIPEC:SegNet (a Mask R-CNN architecture (He et al., 2017)) to segment animal instances in the images or videos first, then it uses SIPEC:PoseNet to estimate the pose for each individual. Later, we can spot a trend of using HRNets and transformers as the 2D APE backbone architecture in step ii) to achieve better performance, especially with the limited animal pose data. For example, the top-down DepthFormer method (Liu et al., 2022b) has been proposed to achieve lightweight multiple-animal pose estimation by designing a new basic block based on the transformer structure and the similarities between self-attention and depthwise convolution. In 2024, SuperAnimal-TopViewMouse (Shaokai Ye and Mathis, 2024) adapts Faster R-CNN for enhanced feature extraction to detect animal instances in the given image or videos first, and then also uses the transformer-based AnimalTokenPose models to achieve high zero-shot APE performance in OOD animal datasets. In the same paper, both SuperAnimal-TopViewMouse (Shaokai Ye and Mathis, 2024) and SuperAnimal-Quadruped (Shaokai Ye and Mathis, 2024) models also enable us to use HRNet-w32 (Wang et al., 2020) as the underlying architecture to get the final animal 2D pose. Until now, SuperAnimal (Shaokai Ye and Mathis, 2024) has claimed to be the current state-of-the-art across a wide variety of animal species for both 2D images and videos. More information on how SuperAnimal is adopted to videos is discussed in section 4.2. Another recent top-down approach for 2D endangered species (i.e. lynx) pose estimation with limited data (Straka et al., 2024) also leverages the high-resolution network HRNet-w32 (Wang et al., 2020) architecture in step ii) to get accurate 2D animal pose, as this HRNet-w32 architecture helps to maintain detailed spatial information throughout the network. All of these methodologies prove the trend.
@@
In contrast, bottom-up methods (LABEL:bbb) prioritize predicting all keypoint locations in an image first, then grouping them into individual instances. Initial work in this area includes methods like OpenPose (Cao et al., 2017), which introduced the concept of Part Affinity Fields (PAFs) to associate body parts with individuals coherently. This PAFs idea was further adapted in the DLCRNet and maDLC architecture (Lauer et al., 2022) to predict the location and orientation of animal limbs and link the specific keypoints to the correct individual. OpenPose (Cao et al., 2017) was also used in (Blanco Negrete et al., 2021) to associate detected keypoints to particular individuals by taking advantage of similarities in skeleton structure between humans and monkeys. Further refinement in the bottom-up approach was made by works such as (Farahnakian et al., 2021), which focused on pig body part association using DeepLabCut (Mathis et al., 2018).
@@
Both top-down and bottom-up methods have limitations and features. For example, top-down approaches are generally less sensitive to scale variance of objects compared to bottom-up methods, as they use bounding boxes to identify each individual first, then estimate the joints of the detected body in that box region. However, they often face difficulties in scenarios involving occlusions, as the detection step may fail to identify all instances accurately. On the other hand, the primary shortcoming of the bottom-up approaches often stems from incorrect assembly. To allow the users to compare the two competing approaches directly, researchers have developed both bottom-up and top-down variants for maDLC (Lauer et al., 2022), SuperAnimal-TopViewMouse (Shaokai Ye and Mathis, 2024) and SLEAP (Pereira et al., 2022). The decision between these two methods depends on the specific requirements and constraints of the application.
@@
Before we move on to the next section, it is worth noting that some of the APE methods explained in section 4.1 can perform pose estimation on both images and videos, as can be seen in table 5 in column ”IV”. In section 4.2, we explain which of these video-based 2D APE methods extract pose frame by frame, and which uses the spatiotemporal signals of video data in an end-to-end manner for pose estimation.
4.2. 2D Animal Pose Estimation from Videos
@@
Table 6. RGB-based 2D APE methods from videos. All the symbols and column meanings are defined in table 5, and any additional symbols are defined here: [Uncaptioned image] is the underlying architecture as 3D-CNN, [Uncaptioned image]+[Uncaptioned image] is the use of CNN-based pose module followed by an Optical Flow module, [Uncaptioned image] is Bayesian Ensembling, [Uncaptioned image] is Probabilistic Graphical Model.
@@
1 2021 OptiFlex (Liu et al., 2021) S [Uncaptioned image], {[Uncaptioned image]+[Uncaptioned image]}, {[Uncaptioned image], [Uncaptioned image]}, [Uncaptioned image] Mouse, Fish, Monkey, Fly aPCK error (%)
2 2022 T-LEAP (Russello et al., 2022) S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Cow PCKh@0.2(%)
3 2022 SLEAP (Pereira et al., 2022) S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Flies, Bees, Mice, Gerbils mAP (%), PCK,
OKS, Localization
4 2022 SemiMultiPose (Blau et al., 2022) SM [Uncaptioned image], {[Uncaptioned image]+[Uncaptioned image]}, [Uncaptioned image], [Uncaptioned image] Mouse, HoneyBee, Fly mAP, OKS
5 2024 Lightning Pose (Biderman et al., 2023) SM [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Mouse, Fish CCA, Pixel Error, AUC
6 2024 SuperAnimal (Shaokai Ye and Mathis, 2024) U [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] AP-10K (Yu et al., 2021b) RMSE (pixels), mAP
Animal Pose (Cao et al., 2019) Normalized Error
4.2.1. Post-Processing Methods to Incorporate Temporal Information
Early APE frameworks process each video frame independently to predict pose keypoints, without inherently incorporating temporal features in an end-to-end manner during training. Temporal consistency is typically addressed through optional post-processing methods, such as smoothing or filtering, to improve performance. Some examples that follow this frame-by-frame approach include DeepLabCut (Mathis et al., 2018), DeepPoseKit (Graving et al., 2019), and LEAP (Pereira et al., 2019).
4.2.2. End-to-End Methods to Incorporate Temporal Information

### v2

@@ 4.1.1 Single Animal
2D single-frame single-animal pose estimation aims to accurately estimate the 2D pose of an animal from a single image. These methodologies share a conceptual foundation with HPE [171] and can be broadly classified into two categories: heatmap-based and regression methods. The former predicts the locations of an animal’s body joints using heatmaps for supervision, necessitating a subsequent postprocessing step to identify the final keypoints based on the predicted heatmaps. In contrast, the latter approach directly regresses the keypoint coordinates of the animal’s body joints, employing an end-to-end deep neural network to establish a mapping from the input image to the body keypoints. Figure 3 illustrates the general architecture of 2D single-frame single-animal pose estimation methods.
(a) Heatmap-based Method [122].
(b) Regression Method using Few-Shot Keypoint Detection [83].
Figure 3: The general architectures of 2D single-frame single-animal pose estimation methods.
@@
Heatmap-based frameworks such as [111, 60, 70, 72, 69, 40, 71, 90] have become popular in 2D APE. These approaches are structured around creating a 2D heatmap, interpreted as a probabilistic representation of body part locations within the image. The architecture of a typical heatmap-based scheme includes i) a CNN backbone to generate heatmaps from input images, followed by ii) a post-processing step to finalize the pose estimations based on these predictions (Figure 3(a)). There are some common architectures to input an image and output the corresponding heatmaps in step i); for instance, [60, 69, 71, 90] uses the ResNet [46] and its variants to produce the heatmap; [70] adapts a variant of the HRNet [135] as the backbone; [72] applies the stacked hourglass [102] architectures. Subsequently, in step ii), the predicted heatmaps are then processed to produce precise 2D spatial coordinates for each body joint using techniques such as peak detection [40, 90, 111]. Although heatmap-based approaches are well known for their robustness against occlusions and variations in appearance, their computational complexity has been critiqued for impeding end-to-end training. Nonetheless, DeepPoseKit [40] proposes an efficient end-to-end heatmap-based approach.
@@
Regression-based approaches directly regress the coordinates of joint locations as continuous values, employing CNNs to learn a mapping from the input image to the joint coordinates. The output layer of the network consists of multiple units, each representing the $x$ and $y$ coordinates of a joint, and is trained to minimize the distance between the predicted and ground truth coordinates. Although regression-based approaches are less common than heatmap-based methods, they have shown great promise [83]. One notable work [83] adopts a regression-based approach to estimate keypoint positions directly, bypassing the intermediate heatmap representation (Figure 3(b)). Their Few-shot Keypoint Detection (FSKD) model is designed to detect keypoints on unseen species with minimal annotated data, which is particularly useful in challenging scenarios where the species have not been previously encountered or the available annotated data is limited. Regression-based approaches offer the advantage of direct joint location estimation without additional peak detection steps, which can benefit real-time inference. However, they may face long-range 2D offset regression challenges, which may lead to a notable performance gap compared to heatmap-based methods.
@@
Table 5: RGB-based 2D APE Methods. In the column I,V: I,V is the methods that can perform pose estimation on both an image and videos, whereas, I is the APE methods that do only on an image. Column L indicates the learning types as follows: S is supervised methods, U is unsupervised methods, SF is self-supervised methods, SM is semi-supervised methods. [Uncaptioned image] is CNN architecture, [Uncaptioned image] is the use of the attention mechanism, [Uncaptioned image] is GAN, [Uncaptioned image] is underlying architecture as vision transformer, [Uncaptioned image] is Multi-Scale Domain Adaptation Module (MDAM). [Uncaptioned image] is the pose estimation of the candidate algorithm given in the keypoint pose format. [Uncaptioned image] is single-camera-view, [Uncaptioned image] is multi-camera-views. [Uncaptioned image] is indoor experimental set-up and [Uncaptioned image] is in-the-wild experimental set-up. Symbols inside the {} indicate that both categories are used in that method. The method citation is available in column Year.
@@
1 2018 DeepLabCut I,V S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Mammals RMSE (pixels)
2 2019 LEAP I,V S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Fly, Mouse RMSE (pixels)
3 2019 DeepPoseKit I,V S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], {[Uncaptioned image], [Uncaptioned image]} Zebra (2019) ([Uncaptioned image]); Euclidean Error,
Desert locust, Fly ([Uncaptioned image]) Bayesian Inference 2019
4 2020 Li et al. I S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Fly, Worm, Zebrafish RMSE, PI-RMSE, PCK, AUC,
PI-PCK, PI-AUC
5 2021b Synthetic-to-Real I U [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Horse, TigDog (2017) PCK@0.05 (%)
6 2022 Kim et al. I SM,U [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] {[Uncaptioned image], [Uncaptioned image]} TigDog (2017), PCK@0.05 (%)
Animal Pose (2019),
Synthetic Animal (2020)
7 2023 ScarceNet I S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] TigDog (2017), mAP (%), PCK@0.05 (%)
AP-10K (2021b)
8 2023 D-Gen I S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Animal Pose (2019), OKS, PCK@0.05 (%)
AP-10K (2021b),
Animal Kingdom (2022)
@@
9 2022 FSKD I S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], {[Uncaptioned image], [Uncaptioned image]} Animal Pose (2019) PCK
@@
10 2022 SIPEC I,V S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Primates, Mice Pearson Correlation, RMSE (pixels),
mAP (%), IoU, Dice Coefficient
11 2022 DepthFormer I S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] AP-10K (2021b) OKS
12 2024 Straka et al. I S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Lynx PCK, OKS, AP
13 2024 SuperAnimal- I,V S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Animal Pose (2019), RMSE (pixels), Normalized Error,
Quadruped AP-10K (2021b) mAP
@@
14 2017 OpenPose I,V S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], {[Uncaptioned image], [Uncaptioned image]} COCO (2014), MPII (2014) mAP (%)
15 2021 Negrete et al. I S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] MSCOCO (2014), AP
MacaquePose (2021)
16 2021 Farahnakian et al. I,V S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Pig (2019) Error (pixels), RMSE (pixels)
@@
17 2022 DLCRNet, maDLC I,V S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], {[Uncaptioned image], [Uncaptioned image]} 3 Mice, 2 Marmosets ([Uncaptioned image]); RMSE (pixels), mAP, PCK
1 Mice ([Uncaptioned image])
Bottom-up &
Top-down:
@@ 1 Mice ([Uncaptioned image])
18 2022 SLEAP I,V S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Flies, Bees, Mice, Gerbils mAP (%), PCK, OKS, Localization
@@
19 2024 SuperAnimal- I,V S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Animal Pose (2019), RMSE (pixels), Normalized Error,
TopViewMouse AP-10K (2021b) mAP
4.1.2 Multiple Animals
@@
We have seen significant developments in markless pose estimation in individual animals. However, extending these methodologies to multiple animals poses unique challenges, particularly in studying social behaviours or animals in their natural environments [157, 37]. For example, it is common for the body parts of one animal to be obscured by another in a multi-animal pose estimation. Another challenge arises from scale variation, where animals at different distances from the camera show differences in apparent sizes. Inter-species variation also introduces additional complexity due to differences in body structures and poses. However, given these challenges, 2D multiple-animal pose estimation research has still garnered significant attention due to its potential applications in various domains, such as animal behaviour analysis, wildlife monitoring, and ecological studies. Methodologies developed in this field are categorized into two main categories: top-down methods [88, 78, 134, 160] and bottom-up methods [16, 101, 34, 64] (Figure 4).
(a) Top-Down Approach: SLEAP Method [112] Instances - Parts.
(b) Bottom-Up Approach: SLEAP Method [112] Parts - Instances.
Figure 4: 2D single-frame multiple-animal pose estimation using SLEAP method [112].
@@
Top-down approaches for multiple-animal pose estimation (Figure 4(a)) generally involve a two-step process: first i) detecting each animal in the input image using pre-trained object detectors and then ii) applying 2D single-animal pose estimation techniques to estimate the pose for each detected instance. Common object detectors in step i) include Faster R-CNN [117], the Single Shot MultiBox Detector (SSD) [79], You Only Look Once (YOLO) v4 [37], and RetinaNet [76]; common 2D APE methods in step ii) includes variants of ResNet, HRNets and transformer. In the early work, SIPEC [88] uses SIPEC:SegNet (a Mask R-CNN architecture [45]) to segment animal instances in the images or videos first, then it uses SIPEC:PoseNet to estimate the pose for each individual. Later, we can spot a trend of using HRNets and transformers as the 2D APE backbone architecture in step ii) to achieve better performance, especially with the limited animal pose data. For example, the top-down DepthFormer method [78] has been proposed to achieve lightweight multiple-animal pose estimation by designing a new basic block based on the transformer structure and the similarities between self-attention and depthwise convolution. In 2024, SuperAnimal-TopViewMouse [160] adapts Faster R-CNN for enhanced feature extraction to detect animal instances in the given image or videos first, and then also uses the transformer-based AnimalTokenPose models to achieve high zero-shot APE performance in OOD animal datasets. In the same paper, both SuperAnimal-TopViewMouse [160] and SuperAnimal-Quadruped [160] models also enable us to use HRNet-w32 [142] as the underlying architecture to get the final animal 2D pose. Until now, SuperAnimal [160] has claimed to be the current state-of-the-art across a wide variety of animal species for both 2D images and videos. More information on how SuperAnimal is adopted to videos is discussed in Section 4.2. Another recent top-down approach for 2D endangered species (i.e. lynx) pose estimation with limited data [134] also leverages the high-resolution network HRNet-w32 [142] architecture in step ii) to get accurate 2D animal pose, as this HRNet-w32 architecture helps to maintain detailed spatial information throughout the network. All of these methodologies prove the trend.
@@
In contrast, bottom-up methods (Figure 4(b)) prioritize predicting all keypoint locations in an image first, then grouping them into individual instances. Initial work in this area includes methods like OpenPose [16], which introduced the concept of Part Affinity Fields (PAFs) to associate body parts with individuals coherently. This PAFs idea was further adapted in the DLCRNet and maDLC architecture [64] to predict the location and orientation of animal limbs and link the specific keypoints to the correct individual. OpenPose [16] was also used in [101] to associate detected keypoints to particular individuals by taking advantage of similarities in skeleton structure between humans and monkeys. Further refinement in the bottom-up approach was made by works such as [34], which focused on pig body part association using DeepLabCut [90].
@@
Both top-down and bottom-up methods have limitations and features. For example, top-down approaches are generally less sensitive to scale variance of objects compared to bottom-up methods, as they use bounding boxes to identify each individual first, then estimate the joints of the detected body in that box region. However, they often face difficulties in scenarios involving occlusions, as the detection step may fail to identify all instances accurately. On the other hand, the primary shortcoming of the bottom-up approaches often stems from incorrect assembly. To allow the users to compare the two competing approaches directly, researchers have developed both bottom-up and top-down variants for maDLC [64], SuperAnimal-TopViewMouse [160] and SLEAP [112]. The decision between these two methods depends on the specific requirements and constraints of the application.
@@
Before we move on to the next section, it is worth noting that some of the APE methods explained in Section 4.1 can perform pose estimation on both images and videos, as can be seen in Table 5 in column IV. In Section 4.2, we explain which of these video-based 2D APE methods extract pose frame by frame, and which uses the spatiotemporal signals of video data in an end-to-end manner for pose estimation.
4.2 2D Animal Pose Estimation from Videos
@@
Table 6: RGB-based 2D APE methods from videos. All the symbols and column meanings are defined in Table 5, and any additional symbols are defined here: [Uncaptioned image] is the underlying architecture as 3D-CNN, [Uncaptioned image]+[Uncaptioned image] is the use of CNN-based pose module followed by an Optical Flow module, [Uncaptioned image] is Bayesian Ensembling, [Uncaptioned image] is Probabilistic Graphical Model. The method citation is available in column Year.
@@
1 2021 OptiFlex S [Uncaptioned image], {[Uncaptioned image]+[Uncaptioned image]}, {[Uncaptioned image], [Uncaptioned image]}, [Uncaptioned image] Mouse, Fish, Monkey, Fly aPCK error (%)
2 2022 T-LEAP S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Cow PCKh@0.2(%)
3 2022 SLEAP S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Flies, Bees, Mice, Gerbils mAP (%), PCK, OKS, Localization
4 2022 SemiMultiPose SM [Uncaptioned image], {[Uncaptioned image]+[Uncaptioned image]}, [Uncaptioned image], [Uncaptioned image] Mouse, HoneyBee, Fly mAP, OKS
5 2023 Lightning Pose SM [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Mouse, Fish CCA, Pixel Error, AUC
6 2024 SuperAnimal U [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Animal Pose (2019), RMSE (pixels), Normalized Error,
AP-10K (2021b) mAP
4.2.1 Post-Processing Methods to Incorporate Temporal Information
Early APE frameworks process each video frame independently to predict pose keypoints, without inherently incorporating temporal features in an end-to-end manner during training. Temporal consistency is typically addressed through optional post-processing methods, such as smoothing or filtering, to improve performance. Some examples that follow this frame-by-frame approach include DeepLabCut [90], DeepPoseKit [40], and LEAP [111].
4.2.2 End-to-End Methods to Incorporate Temporal Information